OpenStudy (anonymous):

A clock radio is on sale for 25% off its original retail price of $30. If the delivery charge for the radio is 10% of its purchase price, how much would it cost to be delivered? $8.25 $15 $20.25 $24.75 $25.50

OpenStudy (espex):

\[.75 \times 30 = x\] \[.1 \times x = delivery\]

OpenStudy (anonymous):

i got 2.25

OpenStudy (espex):

Not one of your answers, but that is correct. *shrug*

OpenStudy (anonymous):

@eSpeX thanx anyways, bro

OpenStudy (espex):

Oh, they are including the purchase price. :D

OpenStudy (anonymous):

\[30\times .75\times 1.1\] should do it

OpenStudy (espex):

Well, the cost to be delivered is the cost of the radio, 22.50, plus the delivery fee, 2.25.

OpenStudy (espex):

Initially I read it to ask what the delivery fee was. :)

OpenStudy (anonymous):

the question is phrased incorrectly because how much would it cost to be delivered? means, well, it means how much would it cost to be delivered, not how much is the total price

OpenStudy (anonymous):

@eSpeX you read what it actually asked. it is the question that is incorrect

OpenStudy (espex):

:)

OpenStudy (anonymous):

@satellite73 how should this be solved?

OpenStudy (espex):

@mitchelsewbaran It already is solved. Read the thread again.